OpenStudy (anonymous):

prove the identity: cscx-sinx=cosxcotx remember that in a proof you can only change one side

OpenStudy (anonymous):

(1/Sin) - Sin = Cos*Cos/Sin (1-Sin^2)/Sin = Cos^2/Sin Use the fact that 1=Sin^2+Cos^2 (Sin^2+Cos^2-Sin^2)/Sin = Cos^2/Sin Cos^2/Sin=Cos^2/Sin, qed :)

OpenStudy (anonymous):

oh wait.. well i didn't technically change cosxcotx, but i did lol

OpenStudy (anonymous):

yah you cant change it!

OpenStudy (anonymous):

its easy to do if you can change it

OpenStudy (anonymous):

Sighh stupid arithmetic, it's easy to do if u don't change it as well.. Csc-Sin=Cos*Cot (1/Sin)-Sin=(1-Sin^2)/Sin=Cos^2/Sin=(Cos/Sin)*Cos=Cot*cos

OpenStudy (anonymous):

woah can you please do it the other way its all thrown together Thanks :)

OpenStudy (anonymous):

OMG man, copy+paste works magic, the equality sign separates them all, you're good to go.

OpenStudy (anonymous):

(1/Sin)-Sin= (1-Sin^2)/Sin= Cos^2/Sin= (Cos/Sin)*Cos= Cot*cos

OpenStudy (anonymous):

Im struggling to understand how the -sinx disappears

OpenStudy (anonymous):

It's magic my friend!

OpenStudy (anonymous):

how??

OpenStudy (anonymous):